OpenStudy (anonymous):

PLEASE CAN SOMEONE HELP ME??? Write an equation of a line that is parallel to y = -1/2 x + 3 and passes through the point (8,-4)?

OpenStudy (anonymous):

have you tried point slope form y – y1 = m(x – x1)? So take (8-4) and sub it in with y1 and x1 This will be y-(-4)=m(x-8) Simplify it so it will turn out as y+4=m(x-8) -4 y=m(x-8) -4 y=mx-12 Good Luck!

OpenStudy (anonymous):

Thank u it helps a bunch! Im just not any good at math

OpenStudy (anonymous):

Do u happen to know any thing about parallel and perpendicular? I just dont understand how to tell if it is! Like (-5x + 3y = 2) or (y = -2/3x - 25) figured it couldnt hurt to ask u!